# ST-MambaSync: The Confluence of Mamba Structure and Spatio-Temporal Transformers for Precipitous Traffic Prediction

### v1

@@ 3.3.2 ST-Mamba block and Regression Layer
3.4 Mamba is a kind of Attention in ResNet
Attention can be seen as a kind of linear regression but not used for prediction. Given a set of data $\{(x_{i},y_{i})\}_{i=1}^{N}$ with $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, for a linear model:
$y$ $=x_{i}^{T}\omega+b,$ (26)
for simplicity, set $b=0$ or absorb $b$ into $\omega$. The data matrix can be denoted as:
$X=\begin{bmatrix}x_{1}^{T}\\ \vdots\\ x_{N}^{T}\end{bmatrix}\in\mathbb{R}^{N\times d},\quad\mathbf{y}=\begin{bmatrix}y_{1}\\ \vdots\\ y_{N}\end{bmatrix}.$ (27)
@@ 3.4 Mamba is a kind of Attention in ResNet
For model prediction on the training data with the least squares solution, we can write the formulation as:
@@
With a the given set of data $\{(x_{i},y_{i})\}_{i=1}^{N}$, for find the solution of $\hat{y}_{1}$, it can be written as:
$\hat{y}_{1}$ $=a_{11}y_{1}+a_{12}y_{2}+\dots+a_{1N}y_{N},$ (30)
$a_{ij}$ $=x_{i}^{T}(X^{T}X)^{-1}x_{j},$ (31)
in this case, the $a_{ij}$s are analogous to the attention scores, which are $QK^{T}$. The model prediction $\hat{y}_{1}$ at $x_{1}$ is the ’similarity’ average over all the training values.
For a new data point $x^{\prime}$, the prediction $\hat{y}(x^{\prime})$ by the model is given by:
$\hat{y}(x^{\prime})=x^{\prime}(X^{T}X)^{-1}X^{T}\mathbf{y}.$ (32)
This can be interpreted as the weighted sum of the training data targets, where the weights are the similarities between the new data point and the training data points. Here, $x^{\prime T}(X^{T}X)^{-1}X^{T}$ can be seen as $QK^{T}$, and $\mathbf{y}$ is the $V$.
In the context of commonly used linear attention on data, given data $X\in\mathbb{R}^{N\times d}$, attention can be formulated as:
$\text{attn}(X)=XW_{Q}W_{K}^{T}X^{T}XW_{V}.$ (33)
Now, we move to the discrete update on SSM, which is shown in the following equation (assumed to be Equation 7), treated as the attention score $a_{ij}$, representing the solution or $Q\cdot V$:
$h_{a+1}$ $=e^{A\Delta_{a}}\left(h_{a}+B_{a}u_{a}e^{-A\Delta_{a}}\Delta A\right)$ (34)
$=e^{A\Delta_{a}}h_{a}+B_{a}u_{a}e^{-A\Delta_{a}}\Delta Ae^{A\Delta_{a}}$ (35)
Here, we set $Q_{m}=B_{a}$, $K_{m}=u_{a}$, the weight $W=e^{A\Delta_{a}}$, and the trainable parameter $\beta=\Delta A$. Thus, it can be rewritten as:
$h_{a+1}=Wh+Q_{m}K_{m}W^{-1}\beta.$ (36)
@@
The first term resembles the attention mechanism, capturing the weighted importance of different inputs. The second term incorporates the updated hidden state, structured analogously to a Residual Network (ResNet), where the hidden state is updated by adding a transformed version of itself.
@@ 4.2 Experiment Setup
All experiments are conducted on a machine with the GPU RTX 3090(24GB) CPU 15. The dataset splits for PEMS-BAY, PEMS03, PEMS04, PEMS07, and PEMS08 are configured as follows: PEMS-BAY is segmented into training, validation, and test sets at a ratio of 7:1:2, while PEMS03, PEMS04, PEMS07, and PEMS08 follow a distribution ratio of 6:2:2. We set the embedding dimension ($d_{f}$) to 24 and the dimension of attention ($d_{a}$) to 80. The architecture includes three layers for both spatial and temporal transformers, with a head count of four. We define both the input and forecast horizon at 1 hour, corresponding to ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}H}=Z=12$. For optimization, we employ the Adam optimizer, initiating with a learning rate of 0.001 that gradually decreases, and a batch size set at 16. To enhance training efficiency, an early stopping criterion is implemented, halting the process if validation error fails to improve over 30 consecutive iterations.
@@
Table 2: Performance comparison of models on PEMS datasets.
Model PEMS03 PEMS04 PEMS07 PEMS08
@@
Table 3: Performance on METR-LA and PEMS-BAY.
@@ 4.3 Ablation Study
ST-Mamba Shao et al. (2024) (mamba 1 layer) 13.40 23.20 9.00 0.43 1.18 14
@@
The results as shown as in Table 4, the performance comparison highlights the trade-offs between prediction accuracy and computational efficiency among different models. While multi-layered STAEformer models offer competitive accuracy, they require higher computational resources and longer inference and training times. On the other hand, the ST-Mamba model demonstrates comparable performance with significantly reduced computational requirements, making it a promising solution for efficient traffic flow prediction tasks. Additionally, the ST-MambaSync model, which combines Mamba and attention mechanisms, achieves the lowest prediction errors while maintaining reasonable computational efficiency, indicating its potential for practical deployment in real-world traffic management systems.
5 Discussion

### v3

@@ 3.3.2 ST-Mamba block and Regression Layer
3.4 Analysis of Mamba and Attention
Lemma 1 (Analogy Between Attention and Linear Regression).
Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$ where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, consider a linear regression model defined by $y=x_{i}^{T}\omega$, where $\omega\in\mathbb{R}^{d}$. The least squares solution for the model coefficients $\hat{\omega}$ is expressed as $\hat{\omega}=(X^{T}X)^{-1}X^{T}\mathbf{y}$, enabling predictions $\hat{y}$ on the training data to be given by $\hat{y}=X\hat{\omega}=X(X^{T}X)^{-1}X^{T}\mathbf{y}$. This form of linear prediction can be interpreted as an attention mechanism, where each prediction is a weighted sum of all outputs, with weights determined by the matrix multiplication $X(X^{T}X)^{-1}X^{T}$.
Proof.
For the given data $\{(x_{i},y_{i})\}_{i=1}^{N}$, define the matrix $X$ and vector $\mathbf{y}$ as:
$X=\begin{bmatrix}x_{1}^{T}\\ \vdots\\ x_{N}^{T}\end{bmatrix}\in\mathbb{R}^{N\times d},\quad\mathbf{y}=\begin{bmatrix}y_{1}\\ \vdots\\ y_{N}\end{bmatrix}\in\mathbb{R}^{N}.$ (28)
@@ 3.4 Analysis of Mamba and Attention
Using this solution, the model prediction on the training data is:
@@
where each element $\hat{y}_{i}$ for $i=1,\ldots,N$ can be expressed as:
$\hat{y}_{i}=\sum_{j=1}^{N}a_{ij}y_{j},$ (31)
with $a_{ij}=x_{i}^{T}(X^{T}X)^{-1}x_{j}$. These coefficients $a_{ij}$ are analogous to attention scores, typically denoted as $QK^{T}$ in attention mechanisms. Each $\hat{y}_{i}$ represents a weighted average over all outputs, where the weights are the similarities (or ’attention’) the model pays to each training value based on $x_{i}$.
For a new data point $x^{\prime}$, the prediction $\hat{y}(x^{\prime})$ by the model is:
$\hat{y}(x^{\prime})=x^{\prime}(X^{T}X)^{-1}X^{T}\mathbf{y},$ (32)
which can be interpreted as the weighted sum of the training data targets, where the weights are the similarities between the new data point $x^{\prime}$ and the training data points, akin to $QK^{T}$, with $\mathbf{y}$ analogous to $V$ in attention mechanisms.
Proposition 1 (Extension to Spatial State Models).
In the context of Spatial State Models (SSMs), the discrete update equation can be modeled through dynamics analogous to attention mechanisms. Specifically, for each state update:
$h_{a+1}=e^{A\Delta_{a}}h_{a}+B_{a}u_{a}e^{-A\Delta_{a}}\Delta Ae^{A\Delta_{a}},$
where $Q_{m}=B_{a}$, $K_{m}=u_{a}$, and $W=e^{A\Delta_{a}}$ play roles analogous to query, key, and weight transformations in attention mechanisms. This results in an attention-like score for each input that directly influences the model’s state updates.
Proof.
@@
The first term resembles the attention mechanism, capturing the weighted importance of different inputs. The second term incorporates the updated hidden state, structured analogously to a Residual Network (ResNet), where the hidden state is updated by adding a transformed version of itself. ∎
Corollary 1.
The attention mechanism formalism enables SSMs to perform weighted sums of inputs, akin to how attention in neural networks aggregates information. This approach enhances the ability of SSMs to dynamically adjust to changing inputs and efficiently compute new states, facilitating more robust predictions and state estimations.
3.5 The iteration expression on ST-MambaSync
In ST-MambaSync, the input vector $\mathbf{Z}$ initially passes through an attention layer and then proceeds to an ST-Mamba Layer. The transformer component primarily captures global information, while the integration with the ST-Mamba block, proven to be an attention mechanism combined with ResNet, complements this by focusing on local details. This combination effectively enhances the transformer’s ability to manage both global and local information, significantly improving the model’s overall predictive accuracy and efficiency.
@@ 4.2 Experiment Setup
All experiments were carried out on a machine equipped with an RTX 3090 GPU (24GB) and a 15-core CPU. The data from the PEMS-BAY, PEMS03, PEMS04, PEMS07, and PEMS08 datasets were divided into training, validation, and test sets. PEMS-BAY was split in a 7:1:2 ratio, whereas PEMS03, PEMS04, PEMS07, and PEMS08 were divided using a 6:2:2 ratio. The embedding dimension ($d_{f}$) was set to 24 and the attention dimension ($d_{a}$) to 80. The model architecture includes a single layer for both spatial and temporal transformers, with four heads, and one ST-Mamba layer with the expansion dimension set to 2. Both the input and forecast horizon were defined as 1 hour, equivalent to ${M}=Z=12$. Optimization was performed using the Adam optimizer, starting with a learning rate of 0.001 that gradually decreased, and a batch size of 16. To improve training efficiency, an early stopping mechanism was implemented, ceasing training if the validation error did not improve after 30 consecutive iterations.
@@
4.3 Performance Evaluation
To assess the effectiveness of the ST-MambaSync model, we utilized six real-world datasets varying significantly in complexity and scale. The datasets range from METR-LA, which includes 207 sensors, to PEMS07, encompassing 883 sensors. This selection provides a broad spectrum of urban traffic patterns and sensor network densities, potentially impacting the predictive performance of the model. The most outstanding results across these evaluations are denoted in red to highlight superior performance.
Analysis of Performance Metrics
The results are shown in Table 2, the ST-MambaSync model showcases its performance across four different PeMS datasets with a focus on three key metrics: MAE, RMSE, and MAPE. Particularly notable is its performance on the PEMS08 dataset, where ST-MambaSync achieves the lowest MAE of 13.30 and RMSE of 23.14, along with a MAPE of 8.80%, indicating superior accuracy compared to other models. Although the model performs consistently across other datasets, such as PEMS03 and PEMS04 with competitive MAEs and RMSEs, it stands out in PEMS08, suggesting its effectiveness in environments with similar traffic patterns and sensor configurations.
Table 2: Performance comparison of models on PEMS datasets, here we denote N as the number of sensors for each dataset.
Model PEMS03 (N=358) PEMS04 (N=307) PEMS07(N=883) PEMS08(N=170)
@@ 4.3 Performance Evaluation
Detailed Analysis Based on Time Horizons
As the performance results showing in Table 3, the models are evaluated across multiple time horizons (15 minutes, 30 minutes, 60 minutes) on the METR-LA and PEMS-BAY datasets. the ST-MambaSync model demonstrates competitive accuracy, particularly excelling in the PEMS-BAY dataset across all time horizons. The visual comparison helps to underscore differences in performance stability and prediction accuracy between models, reaffirming the strengths of ST-MambaSync in shorter forecasting intervals.
Table 3: Performance on METR-LA and PEMS-BAY, we denote N as the number of sensors for each dataset.
@@ 4.4 Ablation Study
In this subsection, we conduct various ablation studies, including adjustments to the layer configurations in both the attention-based and Mamba-based models. Additionally, we analyze the trade-offs between accuracy and computational efficiency.
4.4.1 Vanellia Ablation Study
@@
ST-SSMs(mamba 3 layer) 13.45 23.08 8.96 1.07 3.64 42
ST-SSMs(mamba 2 layer) 13.43 23.14 8.95 0.75 2.56 28
ST-SSMs (mamba 1 layer) 13.40 23.20 9.00 0.43 1.18 14
@@
ST-MambaSync (mamba 1 layer & attention 2 layer) 13.37 23.42 8.98 2.87 3.40 33
ST-MambaSync (mamba 2 layer & attention 1 layer) 13.45 24.16 10.98 1.49 2.96 30
We conducted a comparative analysis with the attention-based model (STAEFormer) (Liu et al., 2023), the Mamba-based model (ST-SSMs) (Shao et al., 2024), and our newly proposed hybrid model (ST-MambaSync). Notably, STAEFormer employs three attention layers. To investigate the impact of the number of attention layers on performance, we modified the STAEFormer by varying the number of layers. As illustrated in Table 4, the comparison underscores the trade-offs between prediction accuracy and computational efficiency across the models. STAEFormer, with multiple attention layers, achieves a competitive mean absolute error (MAE) range of 13.49 to 13.77 but requires substantial computational resources (8.84 to 9.16 FLOPS) and exhibits longer inference (1.20s to 3.03s) and training durations (14s to 36s). Conversely, ST-SSMs, which incorporates a single Mamba layer and no attention layers, shows a high accuracy (MAE of 13.40) with significantly lower computational demands (FLOPS of 0.43), making it an effective option for efficient traffic flow prediction. The ST-MambaSync model, combining one Mamba layer with one attention layer, outperforms others by achieving the lowest prediction error (MAE of 13.30) while maintaining reasonable computational efficiency (1.49 FLOPS). Its inference and training times are 2.65s and 29s, respectively, indicating its suitability for practical implementation in real-world traffic management systems. Increasing the number of Mamba or attention layers in ST-MambaSync, however, results in diminished performance compared to configurations with a single layer of each type. Nevertheless, the integrated approach of combining Mamba and attention layers generally surpasses models utilizing only one of these mechanisms.
4.4.2 Trade Off Analysis on Accuracy and Computation
Figure 2: Trade-offs in Model Performance and Computational Efficiency. This bubble chart illustrates the relationship between Mean Absolute Error (MAE) and computational cost (FLOPS) for various predictive models on the PEMS08 dataset. Each bubble’s size represents the total time required for inference and training, highlighting the efficiency trade-offs. We denote “M” as the number of Mamba layer in the model, “A” as the number of attention layers.
We further conduct a trade-off analysis on the attention-based model and the Mamba-based model, which aids in investigating the combined efficiency of our proposed ST-MambaSync. The results of the trade-off analysis on accuracy and computational effectiveness are presented in Figure 2. The analysis begins by examining two aspects: the increase in attention layers and the increase in Mamba layers.
Increase in Attention Layers
In attention-based models, an increase in the number of attention layers leads to higher FLOPS and computational times, although with improved accuracy. For our integrated model, ST-MambaSync, maintaining a single Mamba layer while increasing the number of attention layers results in a doubling of FLOPS and a significant increase in computational time, without a corresponding improvement in accuracy.
Increase in Mamba Layers
For the Mamba-based model, ST-SSMs, an increase in Mamba layers results in higher FLOPS and extended computational times, but surprisingly, accuracy decreases. This observation suggests that a single Mamba layer is optimal for balancing accuracy and computational efficiency in the integrated ST-MambaSync model. To enhance prediction accuracy without significantly increasing computational burden, incorporating a single attention layer is the most effective strategy. The optimal configuration for achieving the highest accuracy with the least computational trade-off is the ST-MambaSync model with one Mamba layer and one attention layer, which meets these conditions satisfactorily.
4.4.3 Temporal Analysis of Traffic Flow Predictions
@@
5 Discussion and Implication
Figure 4: Comparative Analysis of Prediction Results Using PEMS08 Dataset for Sensors 36 and 127.
This case study evaluates the predictive performance of the STAEformer and ST-MambaSync models across 1-hour, 5-hour, and 24-hour intervals. Each model’s architecture is defined by the number of attention layers (denoted as "A#") and the number of Mamba layers ("M#"). For instance, ST-MambaSync is configured with one attention layer and one Mamba layer, while STAEformer utilizes three attention layers, as illustrated in Figure 4.
Predictions for Sensor 36
  • 1 Hour: Both models diverged from the true values, though ST-MambaSync demonstrated closer approximations at certain intervals.
  • 5 Hours: ST-MambaSync provided more consistent and accurate predictions, closely tracking the actual data, unlike the fluctuating results from STAEformer.
  • 24 Hours: ST-MambaSync showed superior long-term predictive consistency, adhering closely to the actual traffic flow patterns.
Predictions for Sensor 127
  • 1 Hour: Both models performed similarly to those for Sensor 36, with ST-MambaSync slightly more accurate at certain points.
  • 5 Hours: Variability was noted in both models, with neither showing a consistent advantage during the initial hours.
  • 24 Hours: ST-MambaSync maintained closer alignment with the true values, indicating its better capability at handling longer-term dynamics.
The analysis indicates that ST-MambaSync tends to provide more accurate and consistent forecasts across all examined intervals for both sensors. It excels particularly in the 24-hour forecasts, suggesting it is more adept at capturing and adapting to longer-term traffic flow dynamics. ST-MambaSync consistently outperforms the state of art model (SOTA) STAEformer, especially in longer forecast intervals. This emphasizes the importance of selecting appropriate models based on the predictive timeframe and desired accuracy level for traffic management applications.